# COADVISE: Covariate Adjustment with Variable Selection in Randomized Controlled Trials

### v3

@@ 1.1 Background
Despite these regulatory endorsements, significant challenges remain due to high-dimensional covariates and nonlinear relationships. These complexities necessitate robust and comprehensive analytical strategies. However, the comparative advantages and limitations of existing methods remain insufficiently explored, leaving practitioners with limited guidance and uncertainty in selecting the most appropriate approaches. For instance, in the case study presented in Section 5, the Best Apnea Interventions for Research (BestAIR) trial includes data from $N=169$ participants and $p=114$ covariates, making $p$ comparable to $N$. When adjusting for covariates using data from a single treatment arm—where $N_{1}=83$ (treated) and $N_{0}=86$ (controls)—$p$ exceeds the sample sizes. This high-dimensional setting can lead to biased and imprecise covariate-adjusted average treatment effect (ATE) estimates, as the sample size may be insufficient for reliable parameter estimation. Nevertheless, prior studies, such as Gao et al. (2024b), have identified baseline covariate imbalances in the BestAIR trial data, underscoring the need for efficient covariate-adjusted methods. This motivates our approach to improving estimation precision while addressing challenges aforementioned.
@@ 1.2 Related work
Beyond ANCOVA and ANHECOVA, a prominent method in the analysis of RCTs is augmented inverse probability weighting (AIPW) (Ye et al., 2023b, a; Robins et al., 1994; Tsiatis, 2006; Glynn and Quinn, 2010; Cao et al., 2009). The AIPW estimator combines outcome modeling with inverse probability weighting, effectively leveraging information from both observed outcomes and treatment assignment probabilities. This dual-model approach enables AIPW to achieve semiparametric efficiency, which means it attains the semiparametric variance lower bound when the outcome models are correctly specified. As a result, AIPW offers the highest asymptotic efficiency among regular ATE estimators in RCTs (Hahn, 1998; Hirano et al., 2003; Kang et al., 2007; Glynn and Quinn, 2010). One of the defining features of AIPW is its robustness to model misspecification in the context of RCTs. The randomization inherent in trial design allows AIPW to consistently estimate the ATE even when the outcome model is incorrectly specified. This robustness, combined with its efficiency, makes AIPW particularly attractive in practice. Moreover, AIPW offers flexibility by accommodating a wide range of parametric and nonparametric models for outcome estimation, making it more adaptable than ANCOVA or ANHECOVA, which depend on fixed linear model structures. Interestingly, when linear regression is used for outcome modeling within the AIPW framework, the AIPW estimator becomes equivalent to ANHECOVA. In this sense, ANHECOVA can be viewed as a special case of AIPW (Van Lancker et al., 2023). However, this equivalence also underscores a limitation of ANHECOVA: if the linear model is misspecified, the efficiency gains achieved by ANHECOVA are limited compared to those of AIPW when the latter employs correctly specified—or at least better specified—models. For this reason, AIPW is often regarded as a fully adjusted estimator in the literature, capable of optimizing efficiency without being overly reliant on strict parametric assumptions (Li et al., 2023).
@@ 1.3 Our contributions
First, we extend traditional covariate adjustment methods, such as ANCOVA, ANHECOVA, and AIPW, by incorporating a variety of variable selection techniques, including Lasso, adaptive Lasso, marginal correlation, and preliminary testing into the analysis stage of RCTs. This integration addresses the challenges of high-dimensional covariates, enabling efficient and robust estimation of treatment effects. Unlike existing tools such as RobinCar, our framework explicitly supports variable selection, allowing users to identify covariates most predictive of outcomes, thereby improving estimation efficiency.
Second, since missing data are common in RCTs, we provide several missing data imputation methods in our Coadvise R package for user convenience. These methods allow users to impute missing values in the outcome and/or covariates before applying the COADVISE framework. Our package includes complete-case analysis, multiple imputation by chained equations (mice), random forest-based imputation (missForest), inverse probability weighting (ipw), and the use of missingness indicators (missInd). These methods ensure robustness across a range of assumptions about missing data mechanisms, such as missing completely at random (MCAR) or missing at random (MAR). Notably, while tools like RobinCar focus primarily on covariate adjustment, they lack native support for missing data imputation, distinguishing Coadvise as a more comprehensive solution.
To further highlight the differences between Coadvise and RobinCar, we provide a detailed comparison of their capabilities within the context of covariate adjustment for RCTs. Our framework not only supports multi-valued treatments, missing data imputation, and variable selection but also facilitates interoperability by enabling users to export selected covariates for subsequent analysis in RobinCar or custom code. This flexibility ensures that our framework complements rather than replaces existing tools, promoting seamless integration into diverse analytical workflows.
The following infotable (Table 1) provides an overview of our COADVISE framework and functions provided by our Coadvise R package, including available variable selection methods, missing data imputation methods, and covariate-adjusted estimators, along with references to the relevant literature and the properties of different methods. As a remark, based on empirical evidence in Section 4, we recommend the use of Lasso or adaptive Lasso as the variable selection method in general, and we set Lasso as the default choice in our R package.
@@
Missing data imputation Handles missing covariates Handles missing outcome Assumption
Complete-case analysis (cc) $\checkmark$ $\checkmark$ MCAR
Multiple imputation by chained equations (mice; Van Buuren and Groothuis-Oudshoorn (2011)) $\checkmark$ $\checkmark$ MAR
Random forest (missForest; Stekhoven and Bühlmann (2012)) $\checkmark$ $\checkmark$ MAR
Inverse probability weighting (ipw; Robins et al. (1994)) $\times$ $\checkmark$ MAR
Missingness indicator (missInd; Zhao et al. (2024); Zhao and Ding (2022); Song et al. (2021)) $\checkmark$ $\times$ MAR
Variable selection method Lasso (Tibshirani, 1996), Adaptive Lasso (Zou, 2006), Marginal correlation (Permutt, 1990), Preliminary testing (Senn, 1994; Zhao and Ding, 2023)
@@
Software Multi-valued treatments Missing data imputation Variable selection Covariate-adjusted estimators
Coadvise (ours) $\checkmark$ $\checkmark$ $\checkmark$ ANCOVA, ANHECOVA, AIPW
RobinCar (Ye et al., 2023b) $\checkmark$ $\times$ $\times$ ANCOVA, ANHECOVA, AIPW
  • MCAR: missing completely at random; MAR: missing at random.
@@
Table 1: Overview of the COADVISE Framework: The top two sections include all available methods for imputing missing data and variable selection. The third section summarizes available covariate-adjusted estimators with associated variance estimators and their properties. The bottom section provides a comparison with the RobinCar package.
@@ 2.1 Notation and set-up
We consider a binary treatment, denoted by $A\in\{0,1\}$, where $1$ represents the treated group and 00 represents the control group. Each participant in the sample is characterized by a baseline covariate vector $\mathbf{X}$, such as clinical characteristics. We adopt the Neyman-Rubin potential outcomes framework (Neyman, 1923; Imbens and Rubin, 2015), in which each participant has two potential outcomes, $Y(0)$ and $Y(1)$, representing the potential outcomes under control and treatment, respectively.
@@ 2.2 Causal parameter, identification assumptions, and the unadjusted estimator
where the expectation $\mathbb{E}\{\cdot\}$ is taken over the target population represented by the trial sample. However, we can only observe one outcome for each participant, denoted by $Y$. Given the nature and definition of potential outcomes, we assume that the observed outcome is consistent with the potential outcome under the received treatment, i.e., $Y=Y(A)=AY(1)+(1-A)Y(0)$. We further assume that there is only one version of the treatment $A$, and the potential outcome of each participant neither depends on nor affects the treatment received by others (also known as the stable-unit treatment value assumption [SUTVA]). In addition, we make the following standard identification assumptions for RCT data.
@@
There exist two constants $c_{1}$ and $c_{2}$ such that $10<c_{1}<\pi_{1}=P(A=1)<c_{2}<$ 1 almost surely.
@@ 2.4 Non-linear covariate adjustment
$\mathbf{V}=\begin{pmatrix}v_{11}&v_{12}\\ v_{21}&v_{22}\end{pmatrix},$ (7)
@@
$v_{aa}$ $=P(A=a)^{-1}\mathbb{V}\{Y_{a}-\mu_{a}(\mathbf{X})\}+2\text{Cov}\{Y_{a},\mu_{a}(\mathbf{X})\}-\mathbb{V}\{\mu_{a}(\mathbf{X})\},$
@@
The asymptotic variance for $\widehat{\tau}_{\text{AIPW}}=\widehat{\theta}_{1}-\widehat{\theta}_{0}=c^{\prime}\widehat{\boldsymbol{\theta}}:=f(\widehat{\boldsymbol{\theta}})$ is given by $\{\nabla f(\boldsymbol{\theta})\}^{\prime}\mathbf{V}\{\nabla f(\boldsymbol{\theta})\}=v_{11}-2v_{12}+v_{22}$, where $c=(1\ \ -1)^{\prime}$. Therefore, a consistent variance estimator is:
$N^{-1}\{\nabla f(\widehat{\boldsymbol{\theta}})\}^{\prime}\widehat{\mathbf{V}}\{\nabla f(\widehat{\boldsymbol{\theta}})\}=N^{-1}\{\widehat{v}_{11}-2\widehat{v}_{12}+\widehat{v}_{22}\},$
where $\widehat{v}_{11}$, $\widehat{v}_{12}$, and $\widehat{v}_{22}$ are sample-based estimates of $v_{11}$, $v_{12}$, and $v_{22}$, respectively, using the sample versions of the variances, covariances, and proportions.
@@ 3.1 The COADVISE workflow
Once these models are estimated, each model is applied to the data from the opposite treatment group to predict the missing potential outcomes. In other words, a model $\widehat{\mu}_{a}(\boldsymbol{X})$ fitted using data from group $A=a$ is used to predict the potential outcome $Y(a)$ for subjects in group $A=1-a$, for $a=0,1$. This procedure yields all the information necessary to compute the AIPW estimator in (5). Finally, we quantify uncertainty by employing the robust variance estimator in (6) (Ye et al., 2023a).

### v4

@@ 1.1 Background
[…]
@@ 1.2 Related work
Beyond ANCOVA and ANHECOVA, a prominent method in the analysis of RCTs is augmented inverse probability weighting (AIPW) (Robins et al., 1994; Tsiatis, 2006; Glynn and Quinn, 2010; Cao et al., 2009; Kang et al., 2007). The AIPW estimator combines outcome modeling with inverse treatment probability weighting, effectively leveraging information from both observed outcomes and treatment assignment probabilities. This dual-model approach allows AIPW to achieve semiparametric efficiency—that is, to attain the semiparametric variance lower bound—when both the treatment and outcome models are correctly specified. In the setting of RCTs considered in this paper, the treatment assignment mechanism is known by design, so only correct specification of the outcome model is required. As a result, AIPW can be asymptotically most efficient among regular estimators of the ATE in this context. Moreover, AIPW remains consistent even when the outcome model is misspecified, owing to the robustness conferred by randomization. This robustness, combined with its efficiency, makes AIPW particularly attractive in practice. Recent work has extended AIPW methodology to accommodate more complex randomization schemes. For example, Bannick et al. (2023) consider covariate-adaptive randomization, in which treatment assignment probabilities are stratified and potentially correlated with baseline covariates. In certain structured settings, such as when allocation probabilities are constant within strata, AIPW can still achieve semiparametric efficiency with only a correctly specified outcome model. However, under more general covariate-adaptive designs, correct modeling of the treatment assignment mechanism becomes necessary, mirroring challenges found in observational studies. Additionally, Ye et al. (2023a, b) propose robust variance estimators for AIPW under multi-arm RCT settings, which we adopt in our implementation (see Section 2.4).
Moreover, AIPW offers flexibility by accommodating a wide range of parametric and nonparametric models for outcome estimation, making it more adaptable than ANCOVA or ANHECOVA, which depend on fixed linear model structures. Interestingly, when linear regression is used for outcome modeling within the AIPW framework, the AIPW estimator becomes equivalent to ANHECOVA in RCTs considered in this paper. In this sense, ANHECOVA can be viewed as a special case of AIPW (Van Lancker et al., 2023). However, this equivalence also underscores a limitation of ANHECOVA: if the linear model is misspecified, the efficiency gains achieved by ANHECOVA are limited compared to those of AIPW when the latter employs correctly specified—or at least better specified—models. For this reason, AIPW is also regarded as a fully adjusted estimator in the literature, capable of optimizing efficiency without being overly reliant on strict parametric assumptions (Li et al., 2023).
@@ 1.3 Our contributions
First, under a complete randomization scheme, we extend traditional covariate adjustment methods, such as ANCOVA, ANHECOVA, and AIPW, by incorporating a variety of variable selection techniques, including Lasso, adaptive Lasso, marginal correlation, and preliminary testing into the analysis stage of RCTs. This integration addresses the challenges of high-dimensional covariates, enabling efficient and robust estimation of treatment effects. Furthermore, due to concerns regarding the validity of post-selection inference arising from using the same dataset for both variable selection and covariate adjustment (Van Lancker et al., 2024; Belloni et al., 2017), in Section 3.3, we outline regularity conditions that variable selection procedures must satisfy to ensure valid post-selection inference, and we justify that the AIPW estimator remains valid when using Lasso and adaptive Lasso for variable selection.
Second, as detailed in Section 3.4, we introduce two key extensions to the COADVISE framework beyond complete randomization. The first extension incorporates externally trained super-covariates—or outcome foundation models—into the variable selection process. The second extension adapts the framework to covariate-adaptive randomization, allowing treatment allocation to be stratified based on a factor-type variable to achieve improved balance across prognostic strata.
All of the above contributions are implemented in our Coadvise R package. Unlike existing tools such as RobinCar, Coadvise explicitly supports variable selection, enabling users to identify covariates most predictive of outcomes and improve estimation efficiency. To clarify the distinctions between the two packages, we provide a detailed overview of the COADVISE framework and a comparison with RobinCar within the context of covariate adjustment for RCTs (see Table 1). For user convenience, the package includes several missing data imputation methods for handling missing covariates and outcomes, which are common in RCTs; details are provided in Appendix S.1. Our package also promotes interoperability by allowing users to export selected covariates for use in RobinCar or other tools and custom analyses. This flexibility ensures that Coadvise complements rather than replaces existing tools. Based on empirical evidence in Section 4, we recommend using Lasso or adaptive Lasso for variable selection and set Lasso as the default in our package.
@@
Variable selection Lasso (Tibshirani, 1996), Adaptive Lasso (Zou, 2006), Marginal correlation (Permutt, 1990), Preliminary testing (Senn, 1994; Zhao and Ding, 2023)
@@
Software Multi-valued treatments Missing data imputation Variable selection Covariate-adaptive randomization
Coadvise (ours) $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
RobinCar (Ye et al., 2023b) $\checkmark$ $\times$ $\times$ $\checkmark$
@@
Table 1: Overview of the COADVISE Framework. The top two sections summarize all available methods for variable selection and covariate-adjusted estimation, including key properties of the supported estimators. The bottom section compares COADVISE with the RobinCar package. Details on missing data imputation methods are provided in Table S.1 in Appendix S.1.
@@ 2.1 Notation and set-up
We consider a binary treatment, denoted by $A\in\{0,1\}$, where 111 represents the treated group and 0 represents the control group. Each participant in the sample is characterized by a baseline covariate vector $\mathbf{X}$, such as clinical characteristics. We adopt the Neyman-Rubin potential outcomes framework (Neyman, 1923; Imbens and Rubin, 2015), in which each participant has two potential outcomes, $Y(0)$ and $Y(1)$, representing the potential outcomes under control and treatment, respectively.
@@ 2.2 Causal parameter, identification assumptions, and the unadjusted estimator
where the expectation ${}\mathbb{E}\{\cdot\}$ is taken over the target population represented by the trial sample. However, we can only observe one outcome for each participant, denoted by $Y$. Throughout, we assume that (i) the observed outcome is consistent with the potential outcome under the received treatment, i.e., $Y=AY(1)+(1-A)Y(0)$; and (ii) there is only one version of the treatment $A$, and the potential outcome of each participant neither depends on nor affects the treatment received by others (also known as the stable-unit treatment value assumption [SUTVA]). In addition, we make the following standard identification assumptions for RCT data.
@@
There exist a constant $\eta$ such that $0<\eta\leq\pi_{a}=P(A=a)\leq 1-\eta<1$ almost surely, for $a=0,1$.
@@ 2.4 Non-linear covariate adjustment
$\mathbf{V}=\begin{pmatrix}v_{00}&v_{01}\\ v_{10}&v_{11}\end{pmatrix},$ (7)
@@
$v_{aa}$ $=\pi_{a}^{-1}\mathbb{V}\{Y_{a}-\mu_{a}(\mathbf{X})\}+2\text{Cov}\{Y_{a},\mu_{a}(\mathbf{X})\}-\mathbb{V}\{\mu_{a}(\mathbf{X})\},\text{ where }\pi_{a}=P(A=a),$
@@
The asymptotic variance for $\widehat{\tau}_{\text{AIPW}}=\widehat{\theta}_{1}-\widehat{\theta}_{0}=c^{\prime}\widehat{\boldsymbol{\theta}}:=f(\widehat{\boldsymbol{\theta}})$ is given by $\{\nabla f(\boldsymbol{\theta})\}^{\prime}\mathbf{V}\{\nabla f(\boldsymbol{\theta})\}=v_{00}-2v_{01}+v_{11}$, where $c=(1\ \ -1)^{\prime}$. Therefore, a consistent variance estimator is:
$N^{-1}\{\nabla f(\widehat{\boldsymbol{\theta}})\}^{\prime}\widehat{\mathbf{V}}\{\nabla f(\widehat{\boldsymbol{\theta}})\}=N^{-1}\{\widehat{v}_{00}-2\widehat{v}_{01}+\widehat{v}_{11}\},$
where $\widehat{v}_{00}$, $\widehat{v}_{01}$, and $\widehat{v}_{11}$ are sample-based estimates of $v_{00}$, $v_{01}$, and $v_{11}$, respectively, using the sample versions of the variances, covariances, and proportions.
@@ 3.1 The COADVISE workflow
Once these models are estimated, each model is applied to the data from the opposite treatment group to predict the missing potential outcomes. In other words, a model $\widehat{\mu}_{a}(\mathbf{X})$ fitted using data from group $A=a$ is used to predict the potential outcome $Y(a)$ for subjects in group $A=1-a$, for $a=0,1$. This procedure yields all the information necessary to compute the AIPW estimator in (5). Finally, we quantify uncertainty by employing the robust variance estimator in (6) (Ye et al., 2023a).
@@ 3.2 Theoretical results
3.3 Post-selection inference
Although Theorem 1 guarantees efficiency gain regardless of the choice of variable selection method, the uncertainty introduced by variable selection must be accounted for when conducting inference with the final AIPW estimator. This challenge is commonly referred to as the post-selection inference problem. In this section, we outline the theoretical conditions under which a variable selection method enables valid inference using the variance estimator applied to the AIPW estimator constructed from the selected covariates. Design-based theoretical results related to Lasso-based variable selection, in conjunction with the ANHECOVA estimator, have been studied in Bloniarz et al. (2016).
Suppose we have $p>1$ covariates, denoted by $\mathbf{X}=(X_{1},\dots,X_{p})^{\prime}$. Let $\mathcal{S}_{0}\subset\{1,\dots,p\}$ and $\mathcal{S}_{1}\subset\{1,\dots,p\}$ denote the sets of indices corresponding to the relevant covariates for the potential outcomes $Y(0)$ and $Y(1)$, respectively. Define $s_{a}=|\mathcal{S}_{a}|$ for $a=0,1$. We draw upon the theoretical results of Farrell (2015) and Van Lancker et al. (2024) to summarize the following technical conditions within our framework. The former reference considers the more general setting of observational studies, whereas our RCT framework represents a special case that avoids conditions related to the propensity score model. We assume that the estimated outcome regression coefficient $\widehat{\boldsymbol{\beta}}_{a}$ converges in probability to a limiting value $\boldsymbol{\beta}_{a}^{*}$, which may differ from the true coefficient vector $\boldsymbol{\beta}_{a}$ in the outcome model $\mu_{a}(\mathbf{X})=\mu_{a}(\mathbf{X};\boldsymbol{\beta}_{a})$, as model misspecification is allowed in this setting.
Condition 1.
$s_{a}\cdot\log\{\max(p,N)\}/\sqrt{N}=o(1)$ for $a=0,1.$
Condition 2 (Convergence and error rates of estimated coefficients).
$G_{N}^{a}(\mathbf{X})=\frac{1}{N}\sum_{i=1}^{N}\{\mu_{a}(\mathbf{X}_{i};\widehat{\boldsymbol{\beta}}_{a})-\mu_{a}(\mathbf{X}_{i};\boldsymbol{\beta}_{a}^{*})\}^{2}$
is uniformly integrable with $\sqrt{G_{N}^{a}(\mathbf{X})}=o_{P}(1)$, and $\|\widehat{\boldsymbol{\beta}}_{a}-\boldsymbol{\beta}_{a}^{*}\|_{1}=O_{P}\left(s_{a}\sqrt{\log\{\max(p,N)\}/N}\right)$ for $a=0,1$. Here, $P$ is a data generating process that obeys the following regularity and boundedness conditions (below, let $U_{a}=Y(a)-\mu_{a}(\mathbf{X})$ for $a=0,1$):
  • $\{(A_{i},\mathbf{X}_{i},Y_{i})\}_{i=1}^{N}$ is an i.i.d. sample from population $\mathcal{O}=(A,\mathbf{X},Y)$;
  • The covariates $\mathbf{X}$ have bounded support with $\max_{j\in\{1,\dots,p\}}|X_{j}|\leq M<\infty$;
  • $\mathbb{E}\{|U_{a}|^{4}\mid\mathbf{X}\}\leq M<\infty$, $\min_{(j,a)\in\{1,\dots,p\}\times\{0,1\}}\mathbb{E}\{X_{j}^{2}U_{a}^{2}\}$ is bounded away from zero, and for some $r>0$, $\min\{\mathbb{E}\{|\mu_{1}(\mathbf{X})\mu_{0}(\mathbf{X})|^{1+r}\},\mathbb{E}\{|U_{0}|^{4+r}\},\mathbb{E}\{|U_{1}|^{4+r}\}\}\leq M<\infty$, for $a=0,1$.
Condition 1 concerns the relationship among the sparsity $s_{a}$, the number of covariates $p$, and the sample size $N$. This condition is common in the high-dimensional inference literature (Belloni and Chernozhukov, 2013; Belloni et al., 2014, 2017; Bloniarz et al., 2016). Condition 2 requires that the fitted outcome models converge to a bounded limit, along with some regular and mild assumptions on the data-generating process. Both conditions may be strong in more general observational studies, but in our setting, where variable selection is employed and model misspecification is permitted, the dimensionality $p$ can be effectively limited and the existence of the limiting value $\boldsymbol{\beta}_{a}^{*}$ is reasonable.
Outcome models by Lasso and adaptive Lasso variable selections with linear and logistic link functions have been shown to satisfy these conditions (Belloni and Chernozhukov, 2013; Belloni et al., 2017), supporting the validity of post-selection inference using the variance estimator adopted in our framework when either of these two variable selection methods is employed. Lasso also has a standardized implementation in R via the glmnet package, which can be fully pre-specified to prevent data dredging. For user convenience, we have incorporated this function for both Lasso and adaptive Lasso in our Coadvise R package. These methods also address the overfitting issues associated with OLS when the number of covariates is large relative to the sample size.
Finally, we remark on the robustness of variance estimators for covariate-adjusted estimators. Thanks to the RCT design, the propensity score (treatment model) is known and correctly specified. In our framework, variable selection is applied only to the outcome models used in the covariate-adjusted estimators. For all estimators considered—ANCOVA, ANHECOVA, and AIPW with Lasso-based variable selection—Condition 1 ensures that the residuals in their HW sandwich or model-based variance estimators capture the unexplained variation, including uncertainty due to variable selection (Liu et al., 2023; Bloniarz et al., 2016; Zhu et al., 2024). As a result, the robust variance estimators constructed from these residuals implicitly account for this uncertainty.
3.4 Extensions
Thus far, we have discussed our framework under the setting of complete (simple) randomization. In this section, we outline two possible extensions of the COADVISE framework that address emerging challenges in RCT analysis and showcase how our framework may be combined with some advanced techniques. While these extensions are included to illustrate practical implementation possibilities, they are not central contributions of this paper and require more rigorous theoretical investigation in future work. We present them here to help motivate future research directions.
3.4.1 Pre-specified super-covariate
The first extension considers the recently emerging methodology of leveraging super-covariates to enhance RCT analyses. A super-covariate is typically defined or termed as a transformation, summary score, model output, or projection of the existing baseline covariates (potentially high-dimensional) that captures key prognostic information about the outcome, often learned from external data (can be randomized or observational). As noted in several studies (Gagnon-Bartsch et al., 2023; De Bartolomeis et al., 2025), the limited sample sizes in RCTs, together with the increasing availability of large external observational datasets containing the same covariates and outcomes, have motivated growing interest in constructing super-covariates from external data to improve the efficiency of ATE estimation.
Prior research has investigated the super-covariate adjustment approaches. Gagnon-Bartsch et al. (2023) proposed applying machine learning algorithms on multiple external data to obtain foundation models. De Bartolomeis et al. (2025) introduced a hybrid AIPW estimator that integrates available foundation models to improve the efficiency of the AIPW estimator. Schuler et al. (2022); Højbjerre-Frandsen et al. (2025); Liao et al. (2025) and Sun (2024) considered leveraging historical data to learn prognostic regression models. Poulet et al. (2025) extended the idea of prediction powered inference to RCTs, providing externally trained prognostic scores as a function of baseline covariates for all RCT participants.
However, a key concern is that the external data may differ from the RCT data due to distributional shifts, such as covariate shift or outcome shift (Liu et al., 2025; Lee et al., 2023, 2022). As a result, it is generally uncertain whether the super-covariates constructed from external data are truly predictive of the outcomes in the RCT population. To address this issue, we propose a general, data-driven strategy that incorporates both the super-covariates trained from external data and the original covariates in the RCT data for covariate adjustment.
Let $\mathcal{O}=(\mathbf{X},A,Y)$ denote a copy of the RCT data, and in general, suppose we have $K\geq 1$ other mutually independent external data sources for super-covariates. The combination of the COADVISE framework and super-covariates is illustrated in the following Figure 2. We first obtain or train $K$ super-covariates (or models) from each of the $K$ external data sources, denoted by $f_{1}(\mathbf{X}),\dots,f_{K}(\mathbf{X})$. These super-covariates are then applied to the RCT participants to generate corresponding representations, e.g., by predictions. The resulting features are combined with the original covariates $\mathbf{X}$ to form an augmented dataset $\mathcal{O}^{\text{aug}}$, which is subsequently analyzed using the COADVISE framework. Both the original covariates and the super-covariates are treated as candidate variables for selection in the adjustment process.
Figure 2: Illustration of incorporating super-covariates into the COADVISE framework in Figure 1.
In practice, we recommend that users first obtain the $K$ super-covariates and apply them to the RCT data. They can then use our Coadvise R package to perform the analysis based on the augmented RCT dataset, as illustrated in Figure 2. Our proposal offers several flexible features. First, we allow each external dataset to construct super-covariates of arbitrary dimensions, which are then aggregated and incorporated into the augmented dataset as additional covariate information. Second, since COADVISE employs a data-driven variable selection strategy—such as Lasso—even if some of the $K$ super-covariates are not predictive of the outcome in the RCT data, they may be excluded prior to the final adjustment step. This enables COADVISE to automatically guard against the inclusion of irrelevant covariates or super-covariates and to maintain an appropriate balance between sample size and model complexity.
3.4.2 Covariate-adaptive randomization
The preceding discussion in this paper focused on complete (simple) randomization. However, more complex randomization schemes are commonly employed in practice to ensure balance in treatment allocation across key subgroups. For instance, stratified permuted block randomization often assigns participants within predefined blocks (e.g., based on age or clinical site) to treatment arms in a way that maintains a predetermined allocation ratio, thereby achieving better balance within each block (Zelen, 1974).
Our second extension centers around the covariate-adaptive randomization (Bugni et al., 2018; Ye et al., 2022; Rafi, 2023; Bannick et al., 2023) including stratified permuted block randomization (Zelen, 1974) and Pocock and Simon’s minimization (Pocock and Simon, 1975; Taves, 1974) as special cases. In general, we let $\mathbf{Z}$ be a discrete baseline covariate vector used in covariate-adaptive randomization, where $\mathbf{Z}$ has $L<\infty$ distinct joint levels $\mathbf{z}_{1},\dots,\mathbf{z}_{L}$. Also, $\mathbf{Z}$ and $\mathbf{X}$ can share some common components. For example, the gender can be a factor for stratified randomization at the design stage, but it can also be incorporated into the covariate adjustment in the analysis stage. Following Bannick et al. (2023), we impose the following mild assumption on treatment assignment in covariate-adaptive RCTs. We use $\pi_{a}(\mathbf{z})=P(A=a\mid\mathbf{Z}=\mathbf{z})$ to denote the treatment assignment probability for arm $a$ within stratum $\mathbf{z}$.
Assumption 3.
[…]
Assumption 3(ii) ensures that each treatment is assigned with the same marginal probability across all strata defined by $\mathbf{Z}$, which is a primary goal of covariate-adaptive randomization. The estimand of interest remains the ATE, $\tau=\mathbb{E}\{Y(1)-Y(0)\}$. As shown in Bannick et al. (2023), the AIPW estimator (5) remains consistent for $\tau$ even when the outcome model is misspecified. To incorporate covariate-adaptive randomization into the COADVISE framework, we propose using the same AIPW estimation procedure outlined in Figure 1 following variable selection. The only modification lies in the final step, where a generalized variance expression derived by Bannick et al. (2023) is applied.
We first review key aspects of the asymptotic variance expression. Recalling the notation in Equation (6) from Section 2.4, we now present its modification under covariate-adaptive randomization, following Bannick et al. (2023), with stratification by $\mathbf{Z}$:
$\sqrt{N}(\widehat{\boldsymbol{\theta}}-\boldsymbol{\theta})\to_{d}\mathcal{N}(\mathbf{0},\mathbf{V}^{*}),$ (8)
where
$\mathbf{V}^{*}=\mathbf{V}-\mathbb{E}[\mathbf{Q}(\mathbf{Z})\left\{\mathbf{\Omega}_{\text{SR}}-\mathbf{\Omega}(\mathbf{Z})\right\}\mathbf{Q}(\mathbf{Z})],$ (9)
with
$\mathbf{Q}(\mathbf{z})=\text{diag}\left\{\pi_{a}^{-1}\mathbb{E}\{Y(a)-\theta_{a}\mid\mathbf{Z}=\mathbf{z}\}-\pi_{a}^{-1}\mathbb{E}[\mu_{a}(\mathbf{X})-\mathbb{E}\{\mu_{a}(\mathbf{X})\}\mid\mathbf{Z}=\mathbf{z}]\right\},\quad\text{for }a=0,1.$
Furthermore, $\mathbf{\Omega}(\mathbf{z})=\text{diag}\left\{\pi_{0}(\mathbf{z}),\pi_{1}(\mathbf{z})\right\}-\left(\pi_{0}(\mathbf{z})~~\pi_{1}(\mathbf{z})\right)^{\prime}\left(\pi_{0}(\mathbf{z})~~\pi_{1}(\mathbf{z})\right)$ is a $\times 2$ 2 covariance matrix determined by the covariate-adaptive randomization design. In contrast, $\mathbf{\Omega}_{\text{SR}}=\text{diag}\{\pi_{0},\pi_{1}\}-(\pi_{0}~~\pi_{1})^{\prime}(\pi_{0}~~\pi_{1})$ denotes the corresponding matrix under simple randomization.
Under Assumption 3, the covariate-adaptive randomization mechanism assigns equal treatment probabilities within each stratum, implying $\pi_{a}(\mathbf{z})=\pi_{a}$ for all $\mathbf{z}$ and $a$, and thus $\mathbf{\Omega}(\mathbf{Z})=\mathbf{\Omega}_{\text{SR}}$. Consequently, $\mathbb{E}\left\{\mathbf{Q}(\mathbf{Z})(\mathbf{\Omega}_{\text{SR}}-\mathbf{\Omega}(\mathbf{Z}))\mathbf{Q}(\mathbf{Z})\right\}=\mathbf{0}$ and the asymptotic variance still simplifies to $\mathbf{V}^{*}=\mathbf{V}$. Nevertheless, to accommodate potential deviations in finite samples, we estimate and retain this correction term in our variance estimator. Specifically, we estimate the treatment probabilities by $\widehat{\pi}_{a}(\mathbf{z})=N_{a}(\mathbf{z})/N(\mathbf{z})$ and $\widehat{\pi}_{a}=N_{a}/N$ for $a=0,1$. Consistent with the estimator of $\mathbf{V}$ in Equation (7), we use sample-based estimates for all components, including proportions, variances, and covariances, to construct the finite-sample version of the variance estimator. Finally, by the Delta method, the variance estimator for the AIPW estimate of the ATE can be obtained accordingly, following the same derivation presented at the end of Section 2.4.
Additionally, we include two basic estimators without covariate adjustment: the simple unadjusted estimator (1), which remains consistent under Assumption 3, and a stratified unadjusted estimator following Bugni et al. (2018), defined as
$\widehat{\tau}_{\text{strata}}=\sum_{\mathbf{z}\in\mathcal{Z}}\frac{N(\mathbf{z})}{N}\left\{\bar{Y}(1;\mathbf{z})-\bar{Y}(0;\mathbf{z})\right\},$ (10)
where $\mathbf{z}$ indexes the strata defined by $\mathbf{Z}$, $\mathcal{Z}$ denotes the support of $\mathbf{Z}$, $N(\mathbf{z})$ is the number of participants in stratum $\mathbf{z}$, and $\bar{Y}(a;\mathbf{z})$ is the average outcome in treatment arm $a$ within stratum $\mathbf{z}$. A natural plug-in variance estimator for $\widehat{\tau}_{\text{strata}}$ is given by:
$\widehat{\mathbb{V}}\left(\widehat{\tau}_{\text{strata}}\right)=\sum_{\mathbf{z}\in\mathcal{Z}}\left(\frac{N(\mathbf{z})}{N}\right)^{2}\left\{\frac{\widehat{S}_{1}^{2}(\mathbf{z})}{N_{1}(\mathbf{z})}+\frac{\widehat{S}_{0}^{2}(\mathbf{z})}{N_{0}(\mathbf{z})}\right\},$ (11)
where $N_{a}(\mathbf{z})$ is the number of participants receiving treatment $a$ in stratum $\mathbf{z}$, and $\widehat{S}_{a}^{2}(\mathbf{z})$ is the sample variance of outcomes in treatment arm $a$ within stratum $\mathbf{z}$, for $a=0,1$. We provide an R function, CoadviseCAR(), in our Coadvise R package to implement the proposed framework for covariate-adaptive randomized trials, where “CAR” stands for covariate-adaptive randomization. A toy example implementing the above three estimators using Lasso variable selection (for AIPW) is given in Appendix S.5.